OpenStudy (anonymous):

-3(5-15n)=15(3n-2) A) 30 B)-15 C)no solution D) all real numbers

OpenStudy (anonymous):

No solutions

OpenStudy (anonymous):

-3(5-15n)=15(3n-2) -15+45n=45n-30 | + 15 45n = 45n - 30 | - 45n 0 = 30 | no solution